# A Path Signature Framework for Detecting Creative Fatigue in Digital Advertising

### v3

@@ Abstract
This paper introduces a signature-based framework for detecting advertising creative fatigue using path signatures, a geometric representation from rough path theory. Creative fatigue—the degradation of creative effectiveness under repeated exposure—is operationally important in digital marketing because delayed detection can translate directly into avoidable opportunity cost. We reframe fatigue monitoring as a geometric change detection problem: advertising performance trajectories are embedded as paths and represented by truncated (log-)signatures, enabling detection of changes in trend, volatility, and non-linear dynamics beyond simple mean or variance shifts. We further connect statistical detection to managerial decision-making via an explicit quantification of performance loss relative to a benchmark period.
Because proprietary production data cannot be released, we evaluate the proposed framework on a synthetic panel dataset designed to mimic realistic impression volumes and noisy day-to-day CTR dynamics. We define observed CTR as the realised binomial rate $CTR_{t}:=C_{t}/I_{t}$ using daily clicks $C_{t}$ and impressions $I_{t}$. The accompanying CSV also contains a pre-computed CTR field (e.g., due to rounding or upstream derivation), but all modelling and evaluation in this paper use $C_{t}/I_{t}$. Crucially, the dataset does not include injected changepoints; we therefore define an operational ground truth for “fatigue onset” based on a noise-robust CTR estimate and a sustained deterioration relative to a recent-best baseline. We report lead-time (early warning) and alert-burden metrics under this operational definition, and provide a sensitivity analysis over the detector’s primary tuning parameters. The methodology scales linearly in time-series length for fixed signature depth and is suitable for monitoring large creative portfolios.
@@ 1 Introduction
The sustained effectiveness of advertising creative is finite. Over time, as an audience becomes saturated with a particular message or visual, its response diminishes a phenomenon widely known as creative fatigue or ad wear-out. For advertisers managing substantial budgets across numerous campaigns, the timely and accurate identification of this performance degradation is a critical challenge. Failing to recognise the onset of fatigue leads to inefficient allocation of media spend, where budgets continue to support underperforming assets, resulting in a direct and often significant opportunity cost. The core problem, therefore, is not merely acknowledging that fatigue occurs, but developing a systematic and scalable methodology to detect its emergence in a timely manner, enabling data-driven decisions about when to refresh or replace a creative asset.
@@
Traditional approaches to monitoring creative performance have often relied on simple heuristics or basic time-series models. A marketing team might, for example, employ a rule-based system that flags a creative for review if its click-through rate declines for three consecutive days. While intuitive, such methods are prone to false positives, reacting to natural variance rather than a true underlying change in performance. Alternatively, classical time-series models like ARIMA can be employed, but these are often slow to adapt to sudden changes and can be computationally intensive to fit at the scale required by modern advertising platforms, which may serve thousands of individual creative variants simultaneously. Manual analysis by human experts, while valuable, is inherently unscalable and subjective. There is consequently a clear need for a more sophisticated approach that is both mathematically robust and computationally efficient.
We propose a novel methodology for quantifying creative fatigue by applying path signature analysis to the domain of marketing performance measurement. Our central contribution is to reframe the problem from one of simple time-series forecasting to one of geometric feature extraction. We treat the performance trajectory of a creative asset over time as a path in two-dimensional space and compute its path signature. This signature provides a rich, non-parametric summary of the path’s shape, capturing intricate details about its trend, volatility, and oscillatory behaviour. By comparing the signatures of adjacent time windows, we can robustly detect statistically significant changes in the underlying performance pattern. This allows us to pinpoint the onset of creative fatigue with greater precision and reliability than existing methods. We further connect this statistical detection to a tangible business outcome by proposing a framework for quantifying the financial wastage associated with a fatigued creative.
Our contributions are fourfold. First, we introduce a signature distance statistic for monitoring creative performance time series, with an explicit algorithm and complexity analysis suitable for large-scale deployment. Second, we formalise a practical mapping from detected regime changes to an interpretable fatigue signal via segment-level trend classification. Third, we propose a benchmark-relative quantification of performance loss that links detection outputs to managerial prioritisation. Fourth, we provide a technically transparent evaluation protocol for noisy daily CTR data by defining an operational ground truth for fatigue onset and reporting lead-time and alert-burden metrics alongside sensitivity analyses.
@@
The remainder of this paper is structured as follows. We begin by formalising the problem of creative fatigue and introducing the mathematical concept of the path signature. We then present the technical details of our signature-based change point detection methodology, including the algorithm for its implementation. Subsequently, we introduce a framework for quantifying the financial impact of fatigue. The performance of our method is then evaluated through a simulation study on synthetic datasets designed to mirror real-world patterns. Finally, we conclude with a discussion of our findings and suggest avenues for future research.
@@ 2 Literature Review
The challenge of detecting creative fatigue sits at the intersection of multiple research domains: advertising effectiveness theory, time-series analysis, change point detection, and modern computational methods. This comprehensive review synthesises these diverse literatures to establish the theoretical foundation for our methodological contribution and identify the specific gap our signature-based approach addresses.
@@ 2.3 Time-Series Methods in Marketing Science
Pauwels et al. (2004) demonstrate that marketing dynamics often exhibit complex patterns including trends, cycles, and structural breaks, requiring methods that can adapt to multiple types of changes simultaneously. Traditional approaches like ARIMA models, while useful for forecasting, struggle to identify the onset of fatigue due to their assumption of stationarity within windows (Horváth et al., 2014).
The marketing mix modelling literature provides additional context for our work. Ataman et al. (2010) show that the long-term effects of marketing strategies often differ from short-term impacts, necessitating methods that can detect changes across multiple timescales. Srinivasan et al. (2010) further argue for incorporating "mind-set metrics" that capture underlying psychological states—a capability that geometric methods like path signatures naturally provide through their hierarchical feature extraction.
@@ 2.4 Change Point Detection: Methods and Limitations
The statistical literature on change point detection is extensive, with Aminikhanghahi and Cook (2017) and Truong et al. (2020) providing comprehensive surveys of offline and online methods respectively. Classical approaches like CUSUM and MOSUM are computationally efficient but primarily designed for detecting changes in mean or variance, missing the complex dynamic changes that characterise creative fatigue.
Killick et al. (2012) introduce the PELT algorithm for multiple change point detection with linear computational cost, representing a significant advance in scalability. However, their method still assumes piecewise stationarity, which may not hold for advertising data where performance can exhibit gradual degradation or volatile patterns. Adams and MacKay (2007) propose Bayesian online change point detection that can handle more complex patterns, but at significant computational cost that limits scalability to large creative portfolios.
@@ 2.5 Path Signatures: A Geometric Approach to Time-Series
Path signature methods, rooted in rough path theory (Lyons, 1998), offer a fundamentally different approach to time-series analysis. Rather than focusing on statistical moments or model parameters, signatures provide a universal, non-parametric characterisation of a path’s geometry. Chevyrev and Kormilitzin (2016) provide an accessible introduction showing how signatures capture iterated integrals that encode rich information about path behaviour.
The theoretical foundations established by Friz and Hairer (2014) demonstrate that signatures form a graded algebra that uniquely characterises paths up to tree-like equivalences, providing theoretical guarantees absent from many heuristic methods. Recent advances by Kidger et al. (2019) and Morrill et al. (2021) have shown that signature-based methods achieve state-of-the-art performance in various time-series tasks, from finance to healthcare.
Fermanian (2021) establishes that signature embeddings preserve important geometric properties while providing dimensionality reduction, making them computationally tractable for real-time applications. This combination of theoretical rigour and practical efficiency makes signatures particularly suitable for marketing applications where both accuracy and scalability are critical.
@@ 2.6 Research Gap and Our Contribution
First, traditional statistical methods (ARIMA, CUSUM) are designed for simple distributional changes and miss the complex dynamic patterns that characterise creative fatigue. Second, machine learning methods require extensive training data that may not exist for new creatives. Third, current methods lack the scalability required for modern programmatic advertising where thousands of creatives must be monitored simultaneously. Fourth, existing approaches provide binary detection without quantifying financial impact, limiting their utility for budget allocation decisions.
Our contribution addresses these limitations by introducing path signatures to the marketing analytics domain. This geometric approach naturally captures complex temporal patterns without assuming specific functional forms, operates without training data, scales linearly with time-series length, and includes a novel framework for quantifying financial wastage. To our knowledge, this represents the first application of rough path theory to advertising analytics, opening new avenues for geometric methods in marketing science.
The theoretical advantages of signatures—their universal approximation properties, invariance to time reparameterisation, and hierarchical structure—directly address the challenges of creative fatigue detection. By treating performance trajectories as geometric objects rather than statistical processes, we can detect subtle changes that traditional methods miss while maintaining the computational efficiency required for real-time application.
@@ 3 A Formal Framework for Fatigue Detection
where $\otimes$ denotes the tensor product. In our implementation, we compute these signatures using log-signatures (Lie increments) in the free Lie algebra, providing computational advantages while preserving the essential geometric information. The signature provides a non-parametric summary of the path. Chen’s theorem, a cornerstone of rough path theory, guarantees that the signature uniquely determines the path up to tree-like deformations, making it a powerful feature set. Before computing signatures, we apply a normalisation procedure to ensure comparability across different time windows. Specifically, we scale time values to the interval $[0,1]$ and apply min-max scaling to metric values, also mapping them to $[0,1]$. This normalisation ensures that the signature captures the relative shape of the performance trajectory rather than its absolute scale, making the method robust to varying impression volumes and time spans.
Within this framework, we can formally define a change point. We say that a change point $\tau$ occurs in the process $\{Y_{t}\}$ if the law of the process for $t>\tau$ is different from the law of the process for $t\leq\tau$. Our objective is to infer the existence of such a change point from the observed data.
We frame this as a hypothesis testing problem. For any two adjacent time windows, $W_{1}$ and $W_{2}$ of equal size $w_{d}$, we test the null hypothesis $H_{0}$ that the law of the process is the same across both windows, against the alternative hypothesis $H_{1}$ that the law has changed.
$H_{0}:\text{Law}(Y_{t}|_{t\in W_{1}})=\text{Law}(Y_{t}|_{t\in W_{2}})$
$H_{1}:\text{Law}(Y_{t}|_{t\in W_{1}})\neq\text{Law}(Y_{t}|_{t\in W_{2}})$
Our test statistic for this hypothesis is the Euclidean distance between the normalised signatures of the paths corresponding to the two windows:
$D=\|S(X_{W_{1}})-S(X_{W_{2}})\|_{2}$
where $S(X_{W_{i}})$ denotes the truncated signature of the normalised path in window $W_{i}$. We chose the Euclidean distance for its computational efficiency and empirical effectiveness, though other metrics such as the signature kernel distance could be explored in future work. Under the null hypothesis, we expect this distance to be small, reflecting the random variation of a single, consistent process. Under the alternative hypothesis, we expect the distance to be large, reflecting a genuine change in the underlying dynamics of the path. A sufficiently large value of this test statistic leads us to reject the null hypothesis and declare the presence of a change point at the boundary between $W_{1}$ and $W_{2}$.
@@
4 Methodology: Signature-Based Change Point Detection
@@
Our proposed methodology for detecting creative fatigue is based on a systematic algorithmic process designed to identify statistically significant changes in the geometric properties of a performance time series. This section details the algorithm, the selection of its key parameters, and its computational complexity.

### v4

# A Level-Aware Path Signature Framework for Synthetic Screening of Fatigue-Related Advertising Performance Changes

@@ Abstract
This paper introduces a level-aware signature-based framework for screening candidate fatigue-related changes in advertising performance using path signatures, a geometric representation from rough path theory. Creative fatigue—the degradation of creative effectiveness under repeated exposure—is operationally important in digital marketing because delayed review can translate into avoidable opportunity cost. We reframe fatigue monitoring as a geometric change detection problem: advertising performance trajectories are embedded as paths and represented by truncated signatures, providing a structured way to compare adjacent-window trajectory changes. Because fatigue is often expressed as absolute deterioration rather than shape change alone, the primary detector uses calibration-standardised CTR and a directional mean-drop gate. A purely shape-only signature detector is retained as an ablation and generic trajectory-change screen. We further connect statistical screening to operational prioritisation through a benchmark-relative proxy for performance loss.
Because proprietary production data cannot be released, we evaluate the proposed framework on synthetic data designed to mimic plausible impression volumes and noisy day-to-day CTR dynamics. We define observed CTR as the realised binomial rate $CTR_{t}:=C_{t}/I_{t}$ using daily clicks $C_{t}$ and impressions $I_{t}$. The synthetic panel does not include injected changepoints; we therefore define a rule-based evaluation target for “fatigue onset” based on a noise-robust CTR estimate and sustained deterioration relative to a recent-best baseline. We supplement this retrospective target with a stochastic injected-changepoint benchmark containing known fatigue onsets and no-event controls. In that benchmark, the level-aware signature variants recover most injected events and reduce no-event burden relative to CUSUM, GLR-style mean-drop, MOSUM-style mean-drop, and Gaussian-kernel window baselines, while remaining close to a hand-crafted window-feature comparator and behind an offline full-series reference. The methodology scales linearly in time-series length for fixed signature depth, but the present empirical results should be interpreted as synthetic screening evidence rather than validation of online production deployment.
@@ 1 Introduction
The sustained effectiveness of advertising creative is finite. Over time, as an audience becomes saturated with a particular message or visual, its response may diminish, a phenomenon widely known as creative fatigue or ad wear-out. For advertisers managing substantial budgets across numerous campaigns, timely identification of performance changes consistent with fatigue is a critical monitoring challenge. Observed CTR trajectories alone do not identify the cause of deterioration: targeting shifts, auction pressure, placement mix, delivery pacing, and platform learning can produce similar patterns. The core problem addressed here is therefore not causal attribution of fatigue, but systematic and scalable screening for candidate performance changes that warrant analyst review.
@@
Traditional approaches to monitoring creative performance have often relied on simple heuristics or basic time-series models. A marketing team might, for example, employ a rule-based system that flags a creative for review if its click-through rate declines for three consecutive days. While intuitive, such methods are prone to false positives, reacting to natural variance rather than a true underlying change in performance. Alternatively, classical time-series models like ARIMA can be employed, but these are often slow to adapt to sudden changes and can be computationally intensive to fit at the scale required by modern advertising platforms, which may serve thousands of individual creative variants simultaneously. Manual analysis by human experts, while valuable, is inherently unscalable and subjective. There is consequently a need for methods that are transparent, computationally tractable, and able to describe trajectory shape.
We propose a methodology for screening creative fatigue by applying path signature analysis to marketing performance measurement. Our central contribution is to reframe the problem from simple time-series forecasting to geometric feature extraction while retaining fatigue-relevant level information. We treat the performance trajectory of a creative asset over time as a path and compute its path signature. The primary fatigue screen uses calibration-standardised smoothed CTR, adjacent-window signature distances, and a directional mean-drop gate; a shape-only min–max version is evaluated separately as a generic trajectory-change ablation. By comparing adjacent time windows, we flag candidate changes in the observed performance pattern. We further connect this statistical screen to an operational prioritisation metric by proposing a benchmark-relative proxy for the opportunity value associated with underperformance.
Our contributions are fourfold. First, we introduce a level-aware signature distance statistic for monitoring creative performance time series, with an explicit algorithm and complexity analysis. Second, we separate this fatigue-specific detector from a shape-only signature ablation, making clear when signatures are being used as generic trajectory descriptors and when absolute deterioration is retained. Third, we propose a benchmark-relative proxy for performance loss that links screening outputs to managerial prioritisation. Fourth, we provide a synthetic-only evaluation protocol for noisy daily CTR data that combines agreement-with-proxy analysis on a rule-based panel target, an injected-changepoint benchmark with known event times, and a train/held-out threshold-selection check. Detection-accuracy claims are reserved for the injected benchmark; the rule-based panel target is an operational proxy derived from the same CTR trajectories used by the detector.
@@
The remainder of this paper is structured as follows. We begin by formalising the problem of creative fatigue and introducing the mathematical concept of the path signature. We then present the technical details of our signature-based change point detection methodology, including the algorithm for its implementation. Subsequently, we introduce a framework for computing a benchmark-relative impact proxy. The performance of our method is then evaluated through synthetic panel and injected-changepoint studies. Finally, we conclude with a discussion of our findings and suggest avenues for future research.
@@ 2 Literature Review
The challenge of detecting creative fatigue sits at the intersection of advertising effectiveness theory, time-series analysis, change point detection, and modern computational methods. This review summarises the parts of those literatures needed to motivate the signature-based screening framework and its evaluation boundary.
@@ 2.3 Time-Series Methods in Marketing Science
Pauwels et al. (2004) demonstrate that marketing dynamics often exhibit complex patterns including trends, cycles, and structural breaks, requiring methods that can adapt to multiple types of changes simultaneously. Plain ARIMA forecasting models are not designed as online fatigue detectors, although structural-break, intervention, and state-space extensions can address some forms of nonstationarity (Horváth et al., 2014).
The marketing mix modelling literature provides additional context for our work. Ataman et al. (2010) show that the long-term effects of marketing strategies often differ from short-term impacts, necessitating methods that can detect changes across multiple timescales. Srinivasan et al. (2010) further argue for incorporating "mind-set metrics" that capture underlying psychological states. Path signatures do not recover such latent states from CTR alone, but they can encode multiscale temporal structure in observed performance traces that may act as useful proxies for dynamics associated with fatigue.
@@ 2.4 Change Point Detection: Methods and Limitations
The statistical literature on change point detection is extensive, with Aminikhanghahi and Cook (2017) and Truong et al. (2020) providing comprehensive surveys of offline and online methods respectively. Common low-dimensional implementations of CUSUM and MOSUM are computationally efficient and often target changes in means or variances; they can miss richer trajectory dynamics unless the monitored statistic is designed to capture them.
Killick et al. (2012) introduce the PELT pruning algorithm for multiple change point detection with linear computational cost under appropriate conditions, representing a significant advance in scalability. The modelling assumptions depend on the segment cost supplied to PELT; commonly used piecewise-constant or piecewise-stationary costs may be mismatched to advertising data where performance can exhibit gradual degradation or volatile patterns. Adams and MacKay (2007) propose Bayesian online change point detection that can handle more complex patterns, but at significant computational cost that limits scalability to large creative portfolios.
@@ 2.5 Path Signatures: A Geometric Approach to Time-Series
Path signature methods, rooted in rough path theory (Lyons, 1998), offer a fundamentally different approach to time-series analysis. Rather than focusing on statistical moments or model parameters, signatures provide a rich, non-parametric characterisation of a path’s geometry. Chevyrev and Kormilitzin (2016) provide an accessible introduction showing how signatures capture iterated integrals that encode rich information about path behaviour.
The theoretical foundations established by Friz and Hairer (2014) demonstrate that signatures form a graded algebra, while uniqueness results for bounded-variation paths hold up to tree-like equivalence under appropriate hypotheses (Hambly and Lyons, 2010). Recent advances by Kidger et al. (2019) and Morrill et al. (2021) have shown that signature-based methods achieve strong performance in various time-series tasks, from finance to healthcare.
Fermanian (2021) establishes useful properties of signature embeddings for learning from paths. In the present setting, the practical point is that a truncated signature gives a finite-dimensional fixed-depth embedding; it is not inherently a dimensionality reduction unless compared with a richer or longer path representation. This combination of theoretical rigour and fixed-depth tractability makes signatures attractive for marketing applications where auditability and scalability are important.
@@ 2.6 Research Gap and Our Contribution
First, simple traditional monitoring rules and common low-dimensional changepoint statistics are often tuned to specific distributional changes and may miss more complex dynamic patterns that characterise creative fatigue. Second, machine learning methods require extensive training data that may not exist for new creatives. Third, current methods can be difficult to audit at the creative level when many assets must be monitored simultaneously. Fourth, existing approaches often provide binary detection without a transparent prioritisation metric, limiting their utility for budget allocation discussions.
Our contribution addresses these limitations by introducing path signatures to the marketing analytics domain. This geometric approach captures temporal patterns without assuming a specific parametric functional form, can operate without labelled training data, scales linearly with time-series length for fixed signature depth, and includes a transparent framework for computing a benchmark-relative opportunity proxy. We position this as an early domain-specific application of rough path methods to advertising analytics and as a reproducible synthetic benchmark for future comparison.
The theoretical advantages of signatures–their universal approximation properties, reparameterisation properties at the representation level, and hierarchical structure–make them attractive for screening candidate fatigue-related trajectory changes. These properties do not automatically transfer unchanged to the full detector, because the implemented pipeline embeds explicit time, uses fixed observed-day windows, and applies thresholding to finite samples. By treating performance trajectories as geometric objects rather than only as scalar sequences, the approach can detect changes in trajectory shape that simple level-based heuristics may miss, while retaining computational efficiency for fixed signature depth and window size.
@@ 3 A Formal Framework for Fatigue Detection
where $\otimes$ denotes the tensor product. In our implementation, signatures are computed using roughpy over increment streams; log-signature variants are natural alternatives but are not the default evaluated representation in this paper. The signature provides a non-parametric summary of the path. For bounded-variation paths with a fixed starting point, the full signature determines the path up to tree-like equivalence under appropriate hypotheses (Hambly and Lyons, 2010), making it a powerful feature set for path-shaped data. Before computing signatures, we apply a normalisation procedure to improve comparability across different time windows. Specifically, we scale time values to the interval $[0,1]$ and apply min-max scaling to metric values, also mapping them to $[0,1]$. When a metric window is constant, we map the metric coordinate to zero over that window; this makes the signature well-defined and treats flat windows as having no within-window metric movement. This normalisation induces a comparison of within-window relative shape, not the full law or absolute scale of the original process $Y_{t}$. All constant windows are indistinguishable after this transform, and pure level shifts between otherwise flat windows require a separate level feature, global scaling, or an exposure-aware representation. The transform also does not correct for heteroskedastic binomial noise caused by varying impression counts; impression-aware embeddings and exposure-weighted distances are therefore important extensions before production use.
Classical changepoint terminology often defines a change point $\tau$ in a stochastic process $\{Y_{t}\}$ as a time where the law of the process after $\tau$ differs from the law before $\tau$. We use that notion only as background. The implemented detector targets a more local operational object: for a candidate boundary $\tau$, define two adjacent observed-day windows $W_{1}(\tau)=\{\tau-w_{d}+1,\ldots,\tau\}$ and $W_{2}(\tau)=\{\tau+1,\ldots,\tau+w_{d}\}$. A candidate change is flagged when these finite windows have unusually different path-shape summaries. Large local differences can arise from sampling variability as well as true process changes, so the detector should not be read as a test of a global pre/post law-change hypothesis.
Let $\phi_{i}$ denote the window-specific transformation that rescales time to $[0,1]$ and min–max rescales the metric coordinate within window $W_{i}$, using a zero metric coordinate for flat windows. The transformed window path is $\widetilde{X}^{(i)}=\phi_{i}(X_{W_{i}})$. For any two adjacent time windows, $W_{1}$ and $W_{2}$ of equal size $w_{d}$, we compute
$D(\tau)=\|S(\widetilde{X}^{(1)})-S(\widetilde{X}^{(2)})\|_{2},$
where $S(\widetilde{X}^{(i)})$ denotes the truncated signature of the transformed path in window $W_{i}$. This notation is important because signatures are not invariant to arbitrary coordinate scaling; the reported detector compares within-window relative shape rather than raw metric levels. We chose the Euclidean distance for its computational simplicity, though other metrics such as the signature kernel distance could be explored in future work. This score is a heuristic dissimilarity measure rather than a level-controlled statistical test: no null distribution or false-alarm level is claimed for the threshold used below. A sufficiently large value of $D(\tau)$ is treated as a candidate change point at the boundary between $W_{1}$ and $W_{2}$.
@@
4 Methodology: Level-Aware Signature-Based Change Point Detection
@@
Our proposed methodology for screening creative fatigue is based on a systematic algorithmic process designed to identify unusually large changes in the geometric properties of a performance time series while retaining the level deterioration central to fatigue. This section details the primary level-aware detector, the shape-only ablation, the selection of key parameters, and computational complexity.
4.1 Primary level-aware detector
The primary detector used for fatigue screening operates on the same noise-stabilised daily CTR scale used in the injected benchmark. Let $\widehat{p}_{t}$ be the Beta–Binomial posterior CTR estimate and let $\widetilde{CTR}_{t}$ be its trailing moving average. Using an initial calibration period $\mathcal{C}$, define
$z_{t}=\frac{\widetilde{CTR}_{t}-\bar{c}_{\mathcal{C}}}{s_{\mathcal{C}}},$
where $\bar{c}_{\mathcal{C}}$ and $s_{\mathcal{C}}$ are the mean and standard deviation of $\widetilde{CTR}_{t}$ in the calibration period, with a small positive floor on $s_{\mathcal{C}}$ for numerical stability. For each adjacent-window boundary $\tau$, construct two paths from $(u,z_{u})$ over the left and right windows after rescaling local time to $[0,1]$. The level-aware signature distance is
$D_{\mathrm{level}}(\tau)=\|S(X^{L}_{\tau})-S(X^{R}_{\tau})\|_{2}.$
In the past-only version, the threshold at boundary $\tau$ is computed from earlier distances only, after at least $h_{\min}$ earlier distances are available:
$D_{\mathrm{level}}(\tau)>\mu_{<\tau}+k\sigma_{<\tau}.$
A fatigue-specific alert additionally requires a directional deterioration gate,
$\frac{\bar{\widetilde{CTR}}_{L(\tau)}-\bar{\widetilde{CTR}}_{R(\tau)}}{s_{\mathcal{C}}}>\gamma,$
where $\gamma=2$ and $h_{\min}=25$ in the injected benchmark. This gate is not a signature term; it is an explicit domain restriction that distinguishes a fatigue screen from a generic trajectory-change detector. The exposure-aware variant augments the path to $(u,z_{u},\ell_{u})$, where $\ell_{u}$ is calibration-standardised $\log I_{u}$. Scoring boundary $\tau$ requires observing the full right-hand window, so the actionable online alert is available only at the end of that right window, not instantaneously at $\tau$.
Algorithm 1 Past-Only Level-Aware Signature Fatigue Screen
1:Input: Clicks $C_{t}$, impressions $I_{t}$, calibration period $\mathcal{C}$, window size $w_{d}$, depth $d$, threshold multiplier $k$, directional gate $\gamma$, minimum history $h_{\min}$.
2:Compute posterior CTR estimates $\widehat{p}_{t}$ and smoothed series $\widetilde{CTR}_{t}$.
3:Standardise $\widetilde{CTR}_{t}$ to $z_{t}$ using only $\mathcal{C}$.
4:Initialise $Distances\leftarrow[]$ and $Alerts\leftarrow[]$.
5:for each adjacent-window boundary $\tau$ after the calibration period do
6:  Build left and right paths $X^{L}_{\tau},X^{R}_{\tau}$ from $(u,z_{u})$ with local time $u\in[0,1]$.
7:  $D_{\tau}\leftarrow\|S(X^{L}_{\tau})-S(X^{R}_{\tau})\|_{2}$.
8:  if DistancesDistances contains at least $h_{\min}$ past values then
9:   $\theta_{\tau}\leftarrow\mu(D_{<\tau})+k\sigma(D_{<\tau})$.
10:   $G_{\tau}\leftarrow(\bar{\widetilde{CTR}}_{L(\tau)}-\bar{\widetilde{CTR}}_{R(\tau)})/s_{\mathcal{C}}$.
11:   if $D_{\tau}>\theta_{\tau}$ and $G_{\tau}>\gamma$ then
12:     Append boundary $\tau$ and issue time $\tau+w_{d}-1$ to $Alerts$.
13:   end if
14:  end if
15:  Append $D_{\tau}$ to $Distances$.
16:end for
17:return AlertsAlerts.
4.2 Shape-only ablation
The original shape-only detector uses independent min–max scaling in each window. It compares within-window trajectory shape and intentionally discards absolute level and amplitude differences. We retain it because it is useful for understanding what the signature representation alone contributes, but the injected benchmark shows that it is not sufficient for fatigue mechanisms dominated by level deterioration. Consequently, the level-aware detector above is the primary fatigue-screening method evaluated in the synthetic known-onset study.